OpenStudy (anonymous):

Use the quadratic formula to find any x-intercepts on the graph of the equation. Type answer using radicals. y=4x^2-5x-4

OpenStudy (anonymous):

Help....

OpenStudy (anonymous):

Do you know the quadratic formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x-intercept means you let y equal 0 then solve for x. 4x^2-5x-4=0 now use quadratic formula

OpenStudy (anonymous):

so far I have x= -5+square root of 89/8, and I'm stuck

OpenStudy (anonymous):

a = 4, b = -5, c = -4 \[x = \frac{-(-5) \pm \sqrt{(-5)^2 - 4($)(-4)}}{2(4)}\] is that what you got at the beginning?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then it should be \[x = \frac{5\pm \sqrt{89}}{8}\] so 5 should not be negative

OpenStudy (anonymous):

-(-5) = +5

OpenStudy (anonymous):

I do not understand how you got that answer

OpenStudy (anonymous):

I'm stuck on the square root of 89

OpenStudy (anonymous):

i don't think it can be simplified. 89 looks like prime. you can use a calculator to find the exact value of square root of 89 though

OpenStudy (anonymous):

ok, i found that to be true as well, but wanted to double check. The final answer will be x=5+ square root of 89/4?

OpenStudy (anonymous):

no. 5 plus/minus square root of 89/8 in other words \(x = \frac{5 + \sqrt{89}}{8}\) and \(x = \frac{5-\sqrt{89}}{4}\)

OpenStudy (anonymous):

Where are you getting 4 from in the denominator? Should be 8.

OpenStudy (anonymous):

i got confused too. yeah second denominator should be 8